what is the factor given that the size of the population is 500 and samples of size 50 are drawn from this population


1
Expert's answer
2022-02-28T16:58:37-0500

Finite population correction factor: "f=\sqrt{\frac{N-n}{N-1}}=\sqrt{\frac{500-50}{500-1}}\approx0.95."